# Multiscale Quantile Regression with Local Error Control

### v2

@@ Abstract
For robust and efficient detection of change points, we introduce a novel methodology MUSCLE (multiscale quantile segmentation controlling local error) that partitions serial data into multiple segments, each sharing a common quantile. It leverages multiple tests for quantile changes over different scales and locations, and variational estimation. Unlike the often adopted global error control, MUSCLE focuses on local errors defined on individual segments, significantly improving detection power in finding change points. Meanwhile, due to the built-in model complexity penalty, it enjoys the finite sample guarantee that its false discovery rate (or the expected proportion of falsely detected change points) is upper bounded by its unique tuning parameter. Further, we obtain the consistency and the localisation error rates in estimating change points, under mild signal-to-noise-ratio conditions. Both match (up to log factors) the minimax optimality results in the Gaussian setup. All theories hold under the only distributional assumption of serial independence. Incorporating the wavelet tree data structure, we develop an efficient dynamic programming algorithm for computing MUSCLE. Extensive simulations as well as real data applications in electrophysiology and geophysics demonstrate its competitiveness and effectiveness. An implementation via R package muscle is available on GitHub.
@@ 1 Introduction
Detecting and localising abrupt changes in time series data has been a vibrant area of research in statistics and related fields since its early study (e.g. Wald, , 1945; Page, , 1955). See Niu et al., (2016) and Truong et al., (2020) for a comprehensive overview. These methods find applications across a wide range of fields, such as bioinformatics (Futschik et al., , 2014), ecosystem ecology (Berdugo et al., , 2020), electrophysiology (Pein et al., , 2017, 2018), epidemiology (Dehning et al., , 2020), economics and finance (Russell and Rambaccussing, , 2019), molecular biophysics (Fan et al., , 2015) and social sciences (Liu et al., , 2013). The extensive applications have sparked an explosion of methods for detecting changes in the recent literature. Notable examples include the regularised maximum likelihood estimation (e.g. Boysen et al., , 2009; Harchaoui and Lévy-Leduc, , 2010; Killick et al., , 2012; Du et al., , 2016; Maidstone et al., , 2017), the multiscale approaches (e.g. Frick et al., , 2014; Li et al., , 2016; Pein et al., , 2017; Eichinger and Kirch, , 2018) and binary segmentation type methods (e.g. Vostrikova, , 1981; Fryzlewicz, , 2014; Baranowski et al., , 2019; Fang et al., , 2020; Kovács et al., , 2023), to name only a few.
@@ 1.1 Quantile segmentation model
where noises $\varepsilon_{i}$ are independent with their $\beta$-quantiles equal to zero. For instance, the choice of $\beta=0.5$ corresponds to the median segmentation regression. The subscript $\beta$ will be suppressed whenever there is no ambiguity. Since our aim is to provide a $\beta$-quantile segmentation, with an arbitrary yet fixed $\beta\in(0,1)$, we do not consider $\beta$ as a tuning parameter.
We stress that there are no distributional assumptions on observations $Z_{i}$ or equivalently on the additive noises $\varepsilon_{i}$, other than the statistical independence. This allows arbitrarily heavy tailed, skewed or heterogeneous noise distributions, outliers, or a mixture of them (see Section 5.1.1 for some examples). It will further allow $\varepsilon_{i}$ to have a discrete distribution, provided that $\mathbb{P}(\varepsilon_{i}\leq 0)=\beta$. In addition, if we consider multiple quantiles with different $\beta$ in Model 1, we may detect distributional changes in the observations (see Section 5.1.2 and Figure 6).
@@ 1.2 Local error control
and noises $\varepsilon_{i}$ are centred (i.e. $\beta=0.5$) with $\varepsilon_{i}\sim\mathcal{N}(0,1)$ if $1\leq i\leq 100$ and $\varepsilon_{i}\sim t_{3}/(2\sqrt{3})$ if $101\leq i\leq 400$. Here and below $\mathcal{N}(\mu,\sigma^{2})$ denotes the normal distribution with mean $\mu$ and variance $\sigma^{2}$, and $t_{m}$ the Student’s $t$ distribution of $m$ degrees of freedom. To illustrate the impact of data windowing, we apply the proposed MUSCLE method and several recent robust or nonparametric segmentation approaches (Haynes et al., , 2017; Baranowski et al., , 2019; Fearnhead and Rigaill, , 2019; Madrid Padilla et al., , 2021; Vanegas et al., , 2022; Fryzlewicz, , 2024) to two subsets of the data: the first 80 observations and the full set of 400 observations, see Figure 1. All methods, except MUSCLE, produce markedly different estimates (shown as solid and dashed lines) of the signal $f$ on the interval [0, $[0,\,80/n)$ when applied to these two data windows. This reflects the often unspoken fact that the choice of data window acts as a hyper-parameter in real world data applications (cf. Ermshaus et al., , 2023). The reason behind is that many data analysis methods are globally tuned to the specific data window they are applied to.
@@
The distinctive feature of MUSCLE is its stability across different data windows, which primarily stems from its control of local errors on individual segments. Specifically, it controls the probability of falsely detecting change points across all local intervals within every estimated segment. This stability not only simplifies real-data analysis in practical applications but also has important implications for computational efficiency and statistical power. For instance, the data can be split into equally sized segments, MUSCLE can be applied to each segment independently, and the results can then be merged to obtain a final estimate. This splitting-merging approach yields a computation time that scales linearly with the sample size $n$, and enables a linear speedup through parallelisation with multiple computing units (see Section 4 and, in particular, Remark 4). Besides, compared to global error control, local error control, being a weaker criterion, can enhance the power to detect change points, although it may lead to an increased risk of false positives. MUSCLE addresses this by incorporating a built-in penalty on model complexity, with false positive control governed by a single tuning parameter. Further, local error control is especially well suited for change point detection, as a change point is inherently a local feature occurring between two adjacent segments.
@@ 1.3 Our contribution and related work
Under the quantile segmentation model (i.e., Model 1), the proposed MUSCLE scans over intervals of varying scales and locations in search for changes in quantile, and estimates the underlying quantile function in the framework of variational multiscale estimation (Haltmeier et al., , 2022). It penalises the model complexity in terms of the number of change points, under a convex polytope constraint that builds upon a meticulously crafted multiscale statistics towards local error control on each individual segment. Due to this construction, MUSCLE is robust and has a high detection power at the same time, as supported by numerous simulations and real data analyses (see Section 5). Moreover, we showcase two extensions of MUSCLE to the detection of changes in distribution (M-MUSCLE; Section 5.1.2) and to a dependent scenario (D-MUSCLE; Section 5.2.2).
@@
Besides, we have derived consistency (Theorem 3) and bounds on localisation errors (Theorem 4) for MUSCLE in estimating change points, under the detectability condition:
@@
The proposed MUSCLE belongs to the family of multiscale approaches (initiated by Frick et al., , 2014), among which the most closely related methods are FDRSeg (Li et al., , 2016) and MQS (Vanegas et al., , 2022). MUSCLE extends FDRSeg, designed for the standard Gaussian noises, to handle general independent noises, and significantly enhances the detection power of MQS by switching to local error control. Further, the statistical results in this paper are sharper and more general compared to the aforementioned two papers. More precisely, we have improved the upper bound of false discovery rate in Li et al., (2016) from $2\alpha/(1-\alpha)$ to $\alpha$, while also admitting of general distributions beyond Gaussian, by employing a completely different proof technique (see Section 3.2). In comparison with Vanegas et al., (2022), we have refined the overestimation bound on the number of change points (Theorem 2), derived a quantitative estimate of the tail probability of multiscale statistics (Lemma 10 in Appendix A), and reformulated the detection condition into a weaker form (2), which allows frequent large jumps and rare small jumps over long segments to occur at the same time. These improvements are mainly due to the use of strengthened technical inequalities (e.g., Talagrand’s concentration inequality for convex distances) and a subtle adjustment of multiscale statistics (see footnote 1 on page 1). In addition, the optimisation problem defining MUSCLE poses notable challenges compared to FDRSeg and MQS. However, by incorporating the wavelet tree data structure, we are able to achieve nearly (up to a log factor) the same worst-case computational complexity as FDRSeg.
@@
Further related works include robust segmentation methods. Baranowski et al., (2019), though primarily considered Gaussian setups, suggested a heavy tail extension (NOT-HT) that truncates the data into ±$\pm 1$ by a certain sign transform and then treats the truncated data as if they were Gaussian distributed. Fryzlewicz, (2024) further developed this idea, and introduced Robust Narrowest Significance Pursuit (RNSP), which utilises a sign-multiresolution norm. The statistical guarantees of RNSP are provided in the additive noise model (1) with additional assumption that noises are independently and identically distributed (i.i.d.), continuous random variables. From a penalised M-estimation perspective, Fearnhead and Rigaill, (2019) proposed R-FPOP employing Tukey’s biweight loss, and derived asymptotic properties for a restrictive scenario with fixed change points. Another line of related works are nonparametric change point methods, which assume that the observations are piecewise i.i.d. (a stronger model than Model 1; cf. Dümbgen, , 1991). Recently, Matteson and James, (2014) introduced a hierarchical clustering-type method based on the energy distance between probability measures, and proved consistency results (with no rates) under additional moment conditions. Utilising a weighted Kullback–Leibler divergence, Zou et al., (2014) proposed Nonparametric Multiple Change-point Detection (NMCD), together with consistency and localisation rates of change points. As a computational speedup variant of NMCD, Haynes et al., (2017) introduced ED-PELT by simplifying the involved cost function through discrete approximation. Besides, Madrid Padilla et al., (2021) developed a CUSUM procedure (KSD) based on the Kolmogorov–Smirnov distance, and established better statistical guarantees than Zou et al., (2014), but requiring prior knowledge on the minimum segment length. We include a large selection of the above methods in our comparison stduy on simulated and real world data (see Section 5).
@@ 1.4 Outline and notation
Section 2 introduces MUSCLE, a novel methodology for multiple change point segmentation under Model 1. Section 3 delves into the statistical theories of MUSCLE, including consistency, localisation error rates of change points and finite sample guarantees on falsely detected change points. Section 4 discusses the computation of MUSCLE. Section 5 presents simulation studies and real data applications. The paper concludes with a discussion in Section 6. Additional material and proofs are given in the Appendix. The R package muscle implementing MUSCLE and its variants is available on GitHub via https://github.com/liuzhi1993/muscle.
By convention, all intervals in this paper are of form $[a,b)$, with $a,b\in\mathbb{R}$ and $a<b$. For sequences $\{a_{m}\}$ and $\{b_{m}\}$ of positive numbers, we write $a_{m}\lesssim b_{m}$ if $a_{m}\leq Cb_{m}$ for some finite constant $C>0$. If $a_{m}\lesssim b_{m}$ and $b_{m}\lesssim a_{m}$, we write $a_{m}\asymp b_{m}$. We write $X\overset{\mathcal{D}}{=}Y$ to indicate that random variables $X$ and $Y$ have the same distribution. We write $(x)_{+}=\max\{x,0\}$ for $x\in\mathbb{R}$. By $\sqcup$ we denote the disjoint union. By #$\#S$ we denote the cardinality of a set $S$.
@@ 2.1 Multiscale test on a segment
$H_{J}:\mathbb{P}(Z_{i}\leq\theta)=\beta,\quad x_{i}\in J,$
versus $A_{J}:\mathbb{P}(Z_{i}\leq\theta)\neq\beta,\quad x_{i}\in J.$
@@
where $\lvert J\rvert=\#\{i:x_{i}\in J\}$ and $\overline{W}_{J}=\left(\sum_{i\in J}W_{i}\right)/\lvert J\rvert$. Here, we use the convention $00\log 0=$ 0. The null hypothesis $H_{J}$ is rejected if $L_{J}$ exceeds a certain critical value. As in Frick et al., (2014), we incorporate all of such tests with the control of family-wise error, and balance the detection powers over different scales of $\lvert J\rvert$. This leads to the multiscale statistic with scale penalty
@@
which is also referred to as a local quantile on the segment $I$. Note that $\{W_{i}(Z_{i},\theta)\,:\,x_{i}\in I\}$ are i.i.d. Bernoulli distributed with mean $\beta$, under all $H_{J}$ for $J\subseteq I$. Thus, the distribution of $T_{I}(Z,\theta)$ conditioned on all $H_{J}$ for $J\subseteq I$ is known and equal to the maximum of dependent and transformed Bernoulli distributions. A simple and explicit form of $q_{\alpha}(\lvert I\rvert)$ is not yet available, but we can estimate it by Monte–Carlo simulations. Thus, the decision problem of whether $\{Z_{i}:x_{i}\in I\}$ have the same $\beta$-quantile equal to $\theta$ can be addressed by the multiscale test $\mathbbm{1}_{\{T_{I}(Z,\theta)>q_{\alpha}(\lvert I\rvert)\}}$ with the type I error no more than $\alpha$.
@@
The scale penalty term $\sqrt{2\log(e\lvert I\rvert/\lvert J\rvert})$ in (3) is motivated by the observation that $\max_{J:\lvert J\rvert=m}\sqrt{2L_{J}(Z,\theta)}$ asymptotically concentrates around $\sqrt{2\log(\lvert I\rvert/m)}$, see e.g. Kabluchko, (2011). This penalty not only standardises the influence of scale $\lvert J\rvert$ but also ensures the asymptotic tightness of $T_{I}$ as $\lvert I\rvert\to\infty$. The form of scale penalty is though not unique, see e.g.  Walther and Perry, (2022) for alternative forms. Moreover, the distribution of $T_{I}(Z,\theta)$ conditioned on $\cap_{J\subseteq I}H_{J}$ converges weakly to an almost surely finite random variable as $\lvert I\rvert\to\infty$ under a mild restriction on the smallest scale, namely, $\lvert J\rvert\gg\log^{3}\lvert I\rvert$, see e.g. Frick et al., (2014) and Dümbgen and Spokoiny, (2001). It follows that $\limsup_{\lvert I\rvert\to\infty}q_{\alpha}(\lvert I\rvert)<\infty$ for any fixed $\alpha\in(0,1)$. Further, we derive an explicit upper bound of the form $\limsup_{\lvert I\rvert\to\infty}q_{\alpha}(\lvert I\rvert)\lesssim\sqrt{-\log\alpha}$, see Lemma 10 in Appendix A. This bound may serve as an estimate of $q_{\alpha}(\lvert I\rvert)$ particularly when $\lvert I\rvert$ is so large that Monte–Carlo simulations become computationally unaffordable. Alternatively, we might be able to approximate $q_{\alpha}(\lvert I\rvert)$ in a deterministic manner, following ideas in Fang et al., (2020). Their results, however, do not directly apply due to the presence of the scale penalty. The exploration of this possibility is an interesting avenue of future research.

### v3

@@ Abstract
For robust and efficient detection of change points, we introduce a novel methodology MUSCLE (multiscale quantile segmentation controlling local error) that partitions serial data into multiple segments, each sharing a common quantile. It leverages multiple tests for quantile changes over different scales and locations, and variational estimation. Unlike the often adopted global error control, MUSCLE focuses on local errors defined on individual segments, significantly improving detection power in finding change points. Meanwhile, due to the built-in model complexity penalty, it enjoys the finite sample guarantee that its false discovery rate (or the expected proportion of falsely detected change points) is upper bounded by its unique tuning parameter. Further, we obtain the consistency and the localization error rates in estimating change points, under mild signal-to-noise-ratio conditions. Both match (up to log factors) the minimax optimality results in the Gaussian setup. All theories hold under the only distributional assumption of serial independence. Incorporating the wavelet tree data structure, we develop an efficient dynamic programming algorithm for computing MUSCLE. Extensive simulations as well as real data applications in electrophysiology and geophysics demonstrate its competitiveness and effectiveness. An implementation via R package muscle is available on GitHub.
@@ 1 Introduction
Detecting and localizing abrupt changes in time series data has been a vibrant area of research in statistics and related fields since its early study (e.g. Wald, , 1945; Page, , 1955). See Niu et al., (2016) and Truong et al., (2020) for a comprehensive overview. These methods find applications across a wide range of fields, such as bioinformatics (Futschik et al., , 2014), ecosystem ecology (Berdugo et al., , 2020), electrophysiology (Pein et al., , 2018; Requadt et al., , 2025), epidemiology (Dehning et al., , 2020), economics and finance (Russell and Rambaccussing, , 2019), molecular biophysics (Fan et al., , 2015) and social sciences (Liu et al., , 2013). The extensive applications have sparked an explosion of methods for detecting changes in the recent literature. Notable examples include the regularised maximum likelihood estimation (e.g. Boysen et al., , 2009; Harchaoui and Lévy-Leduc, , 2010; Killick et al., , 2012; Du et al., , 2016; Maidstone et al., , 2017), the multiscale approaches (e.g. Frick et al., , 2014; Li et al., , 2016; Pein et al., , 2017; Eichinger and Kirch, , 2018) and binary segmentation type methods (e.g. Vostrikova, , 1981; Fryzlewicz, , 2014; Baranowski et al., , 2019; Fang et al., , 2020; Kovács et al., , 2023), to name only a few.
@@ 1.1 Quantile segmentation model
where noise terms $\varepsilon_{i}$ are independent with their $\beta$-quantiles equal to zero. For instance, the choice of $\beta=0.5$ corresponds to the median segmentation regression. The subscript $\beta$ will be suppressed whenever there is no ambiguity. Since our aim is to provide a $\beta$-quantile segmentation, with an arbitrary yet fixed $\beta\in(0,1)$, we do not consider $\beta$ as a tuning parameter.
We emphasize that no distributional assumptions are imposed on the observations $Z_{i}$, or equivalently on the noise $\varepsilon_{i}$, beyond statistical independence. This framework accommodates arbitrarily heavy-tailed, skewed, or heterogeneous noise distributions, as well as outliers or mixtures thereof (see Section 5.1.1 for examples). Moreover, the noise $\varepsilon_{i}$ may follow discrete distributions, provided that $\mathbb{P}(\varepsilon_{i}\leq 0)=\beta$. When multiple quantile levels $\beta$ are considered in Model 1, the proposed method can further detect distributional changes in the observations (see Section 5.1.2 and Figure 6). Finally, if we focus on a single quantile level, the assumption of independence among the observations $Z_{i}$ can be relaxed to the independence of the truncated random variables $\mathbbm{1}_{\{Z_{i}\leq f(x_{i})\}}\equiv\mathbbm{1}_{\{\varepsilon_{i}\leq 0\}}$.
@@ 1.2 Local error control
and additive errors $\varepsilon_{i}$ are centered (i.e. $\beta=0.5$) with $\varepsilon_{i}\sim\mathcal{N}(0,1)$ if $1\leq i\leq 100$ and $\varepsilon_{i}\sim t_{3}/(2\sqrt{3})$ if $101\leq i\leq 400$. Here and below $\mathcal{N}(\mu,\sigma^{2})$ denotes the normal distribution with mean $\mu$ and variance $\sigma^{2}$, and $t_{m}$ the Student’s $t$ distribution of $m$ degrees of freedom. To illustrate the impact of data windowing, we apply the proposed MUSCLE method and several recent robust or nonparametric segmentation approaches (Haynes et al., , 2017; Baranowski et al., , 2019; Fearnhead and Rigaill, , 2019; Madrid Padilla et al., , 2021; Vanegas et al., , 2022; Fryzlewicz, , 2024) to two subsets of the data: the first 80 observations and the full set of 400 observations, see Figure 1. All methods, except MUSCLE, produce markedly different estimates (shown as solid and dashed lines) of the signal $f$ on the interval [0, 80/n)[0,\,80/n) when applied to these two data windows. This reflects the often unspoken fact that the choice of data window acts as a hyper-parameter in real world data applications (cf. Ermshaus et al., , 2023). The reason behind is that many data analysis methods are globally tuned to the specific data window they are applied to.
@@
The distinctive feature of MUSCLE is its stability across different data windows, which primarily stems from its control of local errors on individual segments. Specifically, it controls the probability of falsely detecting change points across all local intervals within every estimated segment. This stability not only simplifies real-data analysis in practical applications, but also has important implications for computational efficiency and statistical power. For instance, the data can be split into equally sized segments, MUSCLE can be applied to each segment independently, and the results can then be merged to obtain a final estimate. This splitting-merging approach yields a computation time that scales linearly with the sample size $n$, and enables a linear speedup through parallelization with multiple computing units (see Section 4 and, in particular, Remark 4). Besides, compared to global error control, local error control, being a weaker criterion, can enhance the power to detect change points, although it may lead to an increased risk of false positives. MUSCLE addresses this by incorporating a built-in penalty on model complexity, with false positive control governed by a single tuning parameter. Further, local error control is especially well suited for change point detection, as a change point is inherently a local feature occurring between two adjacent segments.
@@ 1.3 Our contribution and related work
Under the quantile segmentation model (i.e., Model 1), the proposed MUSCLE scans over intervals of varying scales and locations in search for changes in quantile, and estimates the underlying quantile function in the framework of variational multiscale estimation (Haltmeier et al., , 2022). It penalizes the model complexity in terms of the number of change points, under a convex polytope constraint that builds upon a meticulously crafted multiscale statistics towards local error control on each individual segment. Due to this construction, MUSCLE is robust and has a high detection power at the same time, as supported by numerous simulations and real data analyses (see Section 5). Moreover, we showcase two extensions of MUSCLE to the detection of changes in distribution (M-MUSCLE; Section 5.1.2) and to a dependent scenario (D-MUSCLE; Section 5.2.2).
@@
Besides, we have derived consistency (Theorem 3) and bounds on localization errors (Theorem 4) for MUSCLE in estimating change points, under the detectability condition:
@@
The proposed MUSCLE belongs to the family of multiscale approaches (initiated by Frick et al., , 2014), among which the most closely related methods are FDRSeg (Li et al., , 2016) and MQS (Vanegas et al., , 2022). MUSCLE extends FDRSeg, designed for the standard Gaussian noise, to handle general independent noise, and significantly enhances the detection power of MQS by switching to local error control. Further, the statistical results in this paper are sharper and more general compared to the aforementioned two papers. More precisely, we have improved the upper bound of false discovery rate in Li et al., (2016) from $2\alpha/(1-\alpha)$ to $\alpha$, while also admitting of general distributions beyond Gaussian, by employing a completely different proof technique (see Section 3.2). In comparison with Vanegas et al., (2022), we have refined the overestimation bound on the number of change points (Theorem 2), derived a quantitative estimate of the tail probability of multiscale statistics (Lemma 10 in Appendix A), and reformulated the detection condition into a weaker form (2), which allows frequent large jumps and rare small jumps over long segments to occur at the same time. These improvements are mainly due to the use of strengthened technical inequalities (e.g., Talagrand’s concentration inequality for convex distances) and a subtle adjustment of multiscale statistics (see footnote 1 on page 1). In addition, the optimization problem defining MUSCLE poses notable challenges compared to FDRSeg and MQS. However, by incorporating the wavelet tree data structure, we are able to achieve nearly (up to a log factor) the same worst-case computational complexity as FDRSeg.
@@
Further related works include robust segmentation methods. Baranowski et al., (2019), though primarily considered Gaussian setups, suggested a heavy tail extension (NOT-HT) that truncates the data into $\pm 1$ by a certain sign transform and then treats the truncated data as if they were Gaussian distributed. Fryzlewicz, (2024) further developed this idea, and introduced Robust Narrowest Significance Pursuit (RNSP), which utilizes a sign-multiresolution norm, and focuses on changes in medians. From a penalized M-estimation perspective, Fearnhead and Rigaill, (2019) proposed R-FPOP employing Tukey’s biweight loss, and derived asymptotic properties for a restrictive scenario with fixed change points. Another line of related works are nonparametric change point methods, which assume that the observations are piecewise i.i.d. (a stronger model than Model 1; cf. Dümbgen, , 1991). Recently, Matteson and James, (2014) introduced a hierarchical clustering-type method based on the energy distance between probability measures, and proved consistency results (with no rates) under additional moment conditions. Utilizing a weighted Kullback–Leibler divergence, Zou et al., (2014) proposed Nonparametric Multiple Change-point Detection (NMCD), together with consistency and localization rates of change points. As a computational speedup variant of NMCD, Haynes et al., (2017) introduced ED-PELT by simplifying the involved cost function through discrete approximation. Besides, Madrid Padilla et al., (2021) developed a CUSUM procedure (KSD) based on the Kolmogorov–Smirnov distance, and established better statistical guarantees than Zou et al., (2014), but requiring prior knowledge on the minimum segment length. We include a large selection of the above methods in our comparison study on simulated and real world data (see Section 5).
@@ 1.4 Outline and notation
Section 2 introduces MUSCLE, a novel methodology for multiple change point segmentation under Model 1. Section 3 delves into the statistical theories of MUSCLE, including consistency, localization error rates of change points and finite sample guarantees on falsely detected change points. Section 4 discusses the computation of MUSCLE. Section 5 presents simulation studies and real data applications. The paper concludes with a discussion in Section 6. Additional material and proofs are given in the Appendix. The R package muscle implementing MUSCLE and its variants is available on GitHub via https://github.com/liuzhi1993/muscle.
[…]
@@ 2.1 Multiscale test on a segment
$H_{J}:\mathbb{P}(Z_{i}\leq\theta)=\beta,\qquad x_{i}\in J,$
versus $A_{J}:\mathbb{P}(Z_{i}\leq\theta)\neq\beta,\qquad x_{i}\in J.$
@@
where $\lvert J\rvert=\#\{i:x_{i}\in J\}$ and $\overline{W}_{J}=\left(\sum_{i\in J}W_{i}\right)/\lvert J\rvert$. Here, we use the convention $0\log 0=0$. The null hypothesis $H_{J}$ is rejected if $L_{J}$ exceeds a certain critical value. As in Frick et al., (2014), we incorporate all of such tests with the control of family-wise error, and balance the detection powers over different scales of $\lvert J\rvert$. This leads to the multiscale statistic with scale penalty
@@
[…]
@@
[…]